Calculate the pH of the solution you will prepare Just part b and c, rates are on left please
Answer #1

All you have to do here, with the graph it's obtain the slope for each of them, this will give you the order of reaction.

r = K[A]n

log r = logK + nlogA

If you do this for the compounds A, B and C in the run, this would be the linear equation you get.

For A:

y = -1.71599 + 1.9221x r2 = 0.999

Which means that the order of reaction for A is 2.

For B:

y = -5.9784 - 0.8383x

The order of B is 1

For C:

y = -4.5519 + 1.0448x

The order of C is 1.

I have a doubt in the last question, but I'll assume that the products will be more dilluted. Hope this helps, and tell me in a comment if you need something else or something to be fixed.